# TuneShield: Mitigating Toxicity in Conversational AI while Fine-tuning on Untrusted Data

### v1

@@ VII-C Adversarial attacks in Dialog-based learning
Multi-level filter [78]  0.16 15.56 1.7 18.08
TuneShield (CH)  0.02 0.04 0.02 0.64
@@
Table IX: RTR for the DD-BART model using DBL in an adaptive attack scenario (adversarial backdoor and indiscriminate attacks) from Weeks et al. [78] using TuneShield.
VIII Conclusion and Future Work
@@
We proposed TuneShield, a defense framework to mitigate toxicity while fine-tuning LLMs on untrusted conversational datasets to create chatbots. TuneShield can be easily integrated into existing fine-tuning pipelines. Key innovations in TuneShield include: (1) Uses highly performant LLM-based toxicity classifier using properties of instruction-tuned, safety-aligned LLMs. (2) TuneShield can mitigate toxicity with minimal impact on conversation quality and also reinforce desirable conversational traits. (3) TuneShield can function reliably even with imperfect or biased toxicity classifiers using a novel model alignment process based on DPO and carefully crafted synthetic conversations. (4) TuneShield exhibits resilience against sophisticated adversarial and jailbreak attacks targeting various stages of the defense pipeline.
@@
We propose the following future work directions: (1) Future work can include using safety-aligned LLMs to detect other evolving toxic language categories [73] and it is worth studying how detection improves as these LLMs evolve. (2) In this work, we focus on mitigating toxicity introduced by fine-tuning on untrusted datasets, without specifically addressing the preservation of the base model’s safety alignment. Future work could explore methods to fine-tune base models while retaining their original safety alignment [57]. We discuss the ethical considerations of this work in Appendix Section -A.
Acknowledgments
@@
This work was supported in part by NSF grant 2231002 and the Commonwealth Cyber Initiative, an investment in the advancement of cyber R&D, innovation, and workforce development. Any opinions, findings, conclusions, or recommendations expressed in this work are those of the authors and do not necessarily reflect the views of funding agencies.

### v2

@@ 7.3. Adversarial attacks in Dialog-based learning
Multi-level filter (Weeks et al., 2023)  0.16 15.56 1.7 18.08
Optimus (CH)  0.02 0.04 0.02 0.64
@@
Table 8. RTR for the DD-BART model using DBL in an adaptive attack scenario (adversarial backdoor and indiscriminate attacks) from Weeks et al. (Weeks et al., 2023) using Optimus.
8. Conclusion and Future Work
@@
We proposed Optimus, a defense framework to mitigate toxicity while fine-tuning LLMs on untrusted conversational datasets to create chatbots. Optimus can be easily integrated into existing fine-tuning pipelines. Key innovations in Optimus include: (1) Uses highly performant LLM-based toxicity classifier using properties of instruction-tuned, safety-aligned LLMs. (2) Optimus can mitigate toxicity with minimal impact on conversation quality and also reinforce desirable conversational traits. (3) Optimus can function reliably even with imperfect or biased toxicity classifiers using a model alignment process based on DPO and carefully crafted synthetic conversations. (4) Optimus exhibits resilience against sophisticated adversarial and jailbreak attacks targeting various stages of the defense pipeline.
@@
We propose the following future work directions: (1) Future work can include using safety-aligned LLMs to detect other evolving toxic language categories (Vishwamitra et al., 2024) and it is worth studying how detection improves as these LLMs evolve. (2) In this work, we focus on mitigating toxicity introduced by fine-tuning on untrusted datasets, without specifically addressing the preservation of the base model’s safety alignment. Future work could explore methods to fine-tune base models while retaining their original safety alignment (Qi et al., 2023).
9. Acknowledgements
@@
This work was supported in part by NSF grants 2231002 and 1943351. Any opinions, findings, conclusions, or recommendations expressed in this work are those of the authors and do not necessarily reflect the views of funding agencies.
Appendix A Appendix
A.1. Ethics Statement
Our evaluation uses only publicly released datasets and locally deployed LLMs, requiring no queries to external services or human subjects. All experiments were conducted in a controlled lab setting. We have not deployed toxicity-injected chatbots and do not intend to release them to prevent potential misuse. Our use of industry API services for toxicity classification was fully compliant with their terms of use. We primarily relied on automated metrics to analyze toxic language, limiting human exposure to a few samples analyzed for false positives and negatives. We believe the research benefits outweigh the risks of this limited exposure.
A.2. Open Science Statement
Our work opens new directions for robust toxicity mitigation. We will release all code and datasets to facilitate future research on securing model customization pipelines.
A.3. Model Fine-tuning and Alignment using DPO
We take our BB, BART and LLaMA-2 models from HuggingFace website (hf_, 2024). BB and BART models are fine-tuned using default settings of AdamW optimizer with a batch size of 64. We fine-tune the BB model for 4 epochs with an LR of 1e-5 and 7e-6 on the offensive and specialized training datasets respectively. Similarly, we fine-tune the BART model with LR of 1e-5 for 10 and 8 epochs respectively on the offensive and specialized datasets. We fine-tune the LLaMA-2 model using QLoRA with 4-bit Normalfloat, bf16 computation datatype and double quantization using the Paged-AdamW optimizer for 3 epochs for both categories. We use a rank $r=$ 64, $\alpha=$ 16, LoRA dropout rate of 10%, LR of 2e-4, batch size per device of 8, gradient accumulation steps of 32 and gradient clip norm of 0.3.
Model alignment using DPO.   We fine-tune the LLaMA-2 model using DPO with same set of hyperparameters but with a LR of 5e-6 and $\beta=0.3$. We fine-tune for 2 and 3 epochs respectively for the offensive and specialized categories.
Adversarial Robustness: Dialog-based Learning.   We use the same hyperparameters from Weeks et al. (Weeks et al., 2023) and fine-tune DD-BART model in the DBL setting. We align the fine-tuned model using DPO for 1 epoch with a batch size of 64, a LR of 5e-7, and gradient accumulation steps of 1 using a paged-AdamW optimizer.
A.4. Toxicity Evaluation Classifiers
We train a BERT-based classifier for each category, to evaluate the RTR for toxic and non-toxic contexts. Table 9 shows the dataset statistics and performance metrics of the toxic class after precision tuning for both classifiers. Both classifiers are trained on imbalanced training datasets using a focal loss (Lin et al., 2017) objective for 5 epochs with an LR of 5e-6, batch size of 16 and an early stopping criteria of 300 steps. We further precision-tune both the toxicity classifiers on the validation sets to ensure higher precision and lower false positives.
Category Class / Source Dataset size
Metrics after
precision tuning
Train Val Precision Recall
  Offensive PersonaChat 40K 4.5K 91.27 59.93
Benign 12K 1.5K
Toxic 12K 1.5K
Specialized PersonaChat 5.5K 0.75K 80.06 59.67
Benign 2.2K 0.3K
Toxic 2.2K 0.3K
Table 9. Dataset statistics and performance metrics (after precision tuning) for toxicity evaluation classifiers for offensive and specialized categories. Benign and Toxic rows are from corresponding toxic category datasets.
Toxicity classifier with CH data distribution.   To cover the CH distribution in FT-Heal, we train evaluation classifiers by adding 30% CH-generated healing samples. After precision tuning, the Offensive classifier achieves 89.72% precision and 59.93% recall on the toxic class, while the specialized classifier achieves 86.54% precision and 60% recall.
Figure 3. Prompt for the Refusal approach.
Figure 4. Instruction for contextual healing.
Model Family Model size   Offensive category  Specialized category
  Precision Recall F1-score FPR  Precision Recall F1-score FPR
[…]
Table 10. Performance of various LLMs using refusal approach and industry APIs for Toxicity classification for toxic class in the offensive and specialized categories. No standard deviation for the Industry APIs due to single trial evaluation.
Defense setting Fine-tuning with healing data (FT-Heal)
Offensive category Specialized category
$B$ BART LLaMA-2 $B$ BART LLaMA-2
No-attack 0.627 0.662 0.606 0.608 0.666 0.596
Attack 0.589 0.647 0.606 0.608 0.682 0.601
Refusal (NH) 0.593 0.632 0.601 0.572 0.666 0.613
Refusal (CH) 0.593 0.679 0.604 0.604 0.714 0.593
O-API (NH) 0.571 0.630 0.607 0.573 0.662 0.599
O-API (CH) 0.585 0.672 0.609 0.599 0.715 0.607
P-API (NH) 0.580 0.638 0.597 0.572 0.667 0.606
P-API (CH) 0.598 0.687 0.607 0.596 0.715 0.607
Table 11. GRADE utility metric for chatbots fine-tuned with Optimus in FT-Heal setting. ”NH” and ”CH” indicate non-contextual and contextual healing, respectively.
Component     Time (Hours)
  Refusal filter     0.38
Healing data generation (CH)     1.25
Fine-tuning + model alignment     2.49
  Fine-tuning w/o Optimus     1.22
Table 12. Computational costs of Optimus components using 2 NVIDIA A100 GPUs. Results are for the specialized category with 22K samples. Healing data generation (CH) is conducted on 2.2K samples using the LLaMA-2-Chat 13B model. We report time taken for fine-tuning LLaMA-2 chatbot.
Optimus
Specialized category
$\beta$ RTR PPL FBD GRD
0.1 7.5 6.05 0.103 0.62
0.2 8 5.95 0.102 0.618
0.3 10.9 5.86 0.102 0.61
Table 13. Trade-off analysis for alignment using DPO for LLaMA-2 chatbot for specialized category using P-API filter and CH approach, while varying $\beta$ values and keeping constant lr=lr=5e-06 and $epochs=3$.
Learning rate  Optimus
  Specialized category
  Epochs RTR PPL FBD GRD
  5e-07  1 21.8 5.63 0.098 0.602
  2 20.4 5.64 0.101 0.602
  3 18.9 5.65 0.099 0.602
5e-06  1 15.7 5.71 0.101 0.611
  2 11.9 5.85 0.101 0.624
  3 7.5 6.05 0.103 0.62
Table 14. Trade-off analysis for alignment using DPO for LLaMA-2 chatbot for specialized category using P-API filter and CH approach, while varying epoch values and keeping $\beta=0.1$ constant for a specific learning rate (5e-6 or 5e-7).
Figure 5. Prompt template to generate adversarial response. For perturbation strategies, refer to PromptAttack (Xu et al., 2023).
Figure 6. Example attack inputs of various jailbreak (JB) attacks. (a) and (b) show example jailbreak attack inputs targeting the refusal classifier. (c) shows an example jailbreak attack input targeting the healing data generation stage.